# Looking To The Horizon: Probing Evolution in the Black Hole Spectrum With GW Catalogs

### v1

@@ Abstract
The population of black holes observed via gravitational waves currently covers the local universe up to a redshift $z\lesssim 1$, for the most massive merging binaries, or $z\lesssim 0.25$ for low-mass BH binaries (BBH). Evolution of the BBH mass spectrum over cosmic time will be a significant probe of formation channels and environments. We demonstrate a reconstruction of the BBH merger rate, allowing for general dependence on binary masses and luminosity distance or redshift and accounting for selection effects, via iterative kernel density estimation (KDE) with optimized multidimensional bandwidths. Performing such reconstructions under a range of detailed assumptions, we see no significant evidence for the evolution of BBH masses with redshift, over the range where detected events are available; at most, a possible trend towards increasing merger rate with redshift for primary masses $m_{1}\gtrsim 50\,M_{\odot}$ is supported. We compare these findings with previous investigations and caution against over-interpreting the current, sparse, data. Significantly upgraded detectors and/or facilities, and longer observing times, are required to harness any correlations of the BBH mass distribution with redshift.
@@ I Introduction
One essential aspect of population analysis is understanding how the masses of BBHs correlate with redshift, which sheds light on their formation environments and underlying astrophysical processes. Recent studies have employed advanced statistical techniques to explore a possible correlation between BBH mass and redshift. In Rinaldi et al. (2024), a hierarchical Gaussian mixture is applied to changing properties of BBH over cosmic history, finding evidence of two distinct groups of black holes with different masses and redshifts, a lack of symmetric binary mergers, and a potential connection between black hole mass and the environment in which they formed. On the other hand Heinzel et al. (2024a) investigates the similar correlation using gravitational wave data with a method that avoids strong assumptions Heinzel et al. (2024b), finding limited evidence for evolution in black hole masses over cosmic time, but possible connections between spin, mass ratio, and distance may exist. Non-parametric population reconstruction has also been proposed to investigate more general questions in GW cosmology and astrophysics in Ng et al. (2024); Fabbri et al. (2025).
@@
Most recently, in Lalleman et al. (2025) the authors focus on the power-law distribution and peak features in the primary mass, carrying out a parameterized change-point analysis to investigate possible changes in such features over redshift: they finding no significant evidence of evolution, although various types of variation in the mass spectrum cannot be excluded.
@@
This study investigates how the mass distribution of BBHs may evolve with redshift using publicly available gravitational wave observations. We employ an improved version of a non-parametric method based on iterative kernel density estimation (KDE), building on previous studies Sadiq et al. (2022, 2024a, 2024b), in combination with an accurate fit of LVK search sensitivity Lorenzo-Medina and Dent (2025). Any evidence for such evolution would provide valuable insights into the astrophysical processes and formation channels that govern BBH systems.
@@
A critical point, given the current small statistics regime of the GW population, is how far statistical inferences and scientific conclusions depend on model assumptions, and conversely, how sensitive they are to fluctuations in the observed GW event set, considered as a random sample. Clearly, little information can be obtained for the population in regions of parameter space where no events are detected. The most obvious statistical limitation is the selection function restricting the visibility of BBH mergers at finite distance, due to detector and search sensitivity: the selection function is also strongly dependent on binary component masses (see e.g. discussion of search sensitivities in Abbott et al. (2023a), and Lorenzo-Medina and Dent (2025)). Thus, the BBH mass spectrum from $\sim\!10\,M_{\odot}$ upward can only be investigated out to fairly small redshift ($z\lesssim 0.25$) with current data.
@@
However even at higher masses, for which the current horizon of detectability extends farther, our knowledge of the BBH population is limited simply by the scarcity of mergers. We already detect a high fraction of high-mass BBH at low redshift, but this is still a very small number of events: thus, only significantly longer observing times will eventually enable an accurate local measurement.
@@
II Methods and data analysis
II.1 Iterative reweighted KDE
@@
Here $\hat{f}_{0}$ is an initial pilot density estimate obtained by setting $\lambda_{i}=1$ and $\alpha$ is the bandwidth sensitivity parameter (lying between 00 and $1$), $g$ being a normalization factor. The optimized choices of the kernel covariance (i.e. bandwidth) and $\alpha$ are determined by maximizing a likelihood figure of merit, evaluated via cross-validation Sadiq et al. (2022). As we use detected events to optimize and evaluate the KDE, reconstruction of the actual astrophysical event density also requires an estimate of the selection function, as discussed below in Sec. III.2.
This method is a computationally efficient approach to obtain reasonable density and rate estimates, along with uncertainties derived by bootstrap resampling. However, the method as previously applied is limited by the use of spherical (isotropic) kernels, implying $\Sigma=h^{2}\mathbb{I}$ in standardized coordinates, where $\mathbb{I}$ represents the identity matrix. This restriction can introduce bias in the estimates when the data follow distinct distributions, or have significantly different measurement uncertainties, over different dimensions.
In this work, we address this limitation by allowing for kernels with different relative bandwidths along each dimension, while the overall kernel scale varies between data points as already described for adaptive KDE. For a data point $\vec{X}_{i}$, the kernel contribution is proportional to
@@
This is not the most general linear data transformation, as we could take $\vec{y}=H^{-1}\vec{x}$ where $H^{-1}$ is any invertible real matrix, corresponding to an off-diagonal kernel bandwidth matrix. We choose not to pursue this more complicated option: such kernels might increase apparent correlations between parameters, thus we expect our conclusions on correlation or evolution to be conservative.
The $D$-dimensional bandwidths $h_{1},\ldots h_{D}$ (acting as rescaling factors) need to be optimized, along with the adaptive parameter $\alpha$. Given the multiple dimensions of this hyper-parameter space, instead of the grid search used in previous works, we switch to a generalized optimization algorithm, in this case chosen as Nelder-Mead. Our cross-validated likelihood figure of merit is not invariant under rescaling coordinates: to ensure stable optimization we evaluate it over the $\vec{x}$ coordinates, i.e. without rescaling  [^1] Standardization also changes the likelihood figure of merit, but only by a constant for a given data set..
Figure 1: Verification of the multidimensional bandwidth optimization using mock data consisting of a mixture of two Gaussian components, represented by red $+$ symbols: contours and blue shading show the KDE in each case. Left: Adaptive KDE with an isotropic kernel (in standardized coordinates): the single bandwidth parameter $h$ and adaptive parameter $\alpha$ optimize the cross-validated likelihood. Right: Adaptive KDE with different optimized bandwidths $h_{1,2}$ for each coordinate.
To validate the multi-dimensional kernel optimization the new adaptive KDE method, we tested it using synthetic data comprising a bimodal Gaussian mixture. This dataset is useful for testing because the isotropic (spherical) kernel lacks the flexibility to capture its multimodality by using different bandwidths per parameter, even after standardizing the data to have the same variance in all dimensions.
We first apply adaptive KDE with an isotropic kernel in standardized coordinates, optimizing a single bandwidth parameter $h$ and adaptive parameter $\alpha$ to maximize the likelihood via leave-one-out cross-validation. However the resulting estimate, shown on the left of Fig. 1, is under-smoothed along one of the two dimensions, introducing excess variance and failing to represent the true distribution. Our improved KDE method allowing independently optimized bandwidths $h_{1}$ and $h_{2}$ for each coordinate better accommodates the anisotropic nature of the data, successfully recovering the bimodal Gaussian mixture, as shown in Fig. 1 (right), as well as significantly increasing the likelihood.
@@ III Application to GWTC-3 data
We apply the reweighted adaptive KDE with multi-dimensional bandwidth optimization to the parameter estimation results LVK (2021a) of gravitational-wave observations cataloged in GWTC-3 Abbott et al. (2023a, c).
@@ III.1 Event selection, parameter estimation and data cleaning
Following the approach in our earlier studies Sadiq et al. (2022, 2024a), we selected 69 BBH candidate events from GWTC-3, which have a false alarm rate below 1 per year. We excluded GW190814, an outlier event with an unusually low secondary mass ($m_{2}\simeq 2.6,M_{\odot}$, $q\simeq 0.11$), as its nature remains uncertain. This system could represent either a very massive neutron star or a light black hole, which is inconsistent with the primary BBH population we aim to analyze Abbott et al. (2020).
In LIGO-Virgo-KAGRA (LVK) catalog parameter estimation analyses, the prior over masses and $d_{L}$ is uniform over redshifted masses $m_{\{1,2\}z}=m_{\{1,2\}}(1+z)$ and over Euclidean volume Abbott et al. (2021a), thus is proportional to $(1+z)^{2}$ $d_{L}^{2}\mathrm{d}m_{1}\mathrm{d}m_{2}\mathrm{d}d_{L}$ (e.g. Callister (2021)). Samples reweighted to correspond to a “cosmological” prior uniform over comoving volume and source-frame time, though still uniform over $m_{z}$, are also available and are the basis of quoted event parameters Abbott et al. (2023a). We use the original samples and apply a factor proportional to the inverse prior $(1+z)^{-2}d_{L}^{-2}$ to our estimate of $p(m_{1}[,m_{2}],d_{L})$ in the reweighting step of our algorithm.
For two events with relatively low SNR, GW190719_215514 and GW190805_211137, we find a small number of PE samples at high distance which appear to represent the prior distribution, which is formally divergent towards large distance, rather than estimating the parameters of the detected merger. The expected (optimal) SNRs for these samples are unusually small. We compare the distances of PE samples with those of simulated signals (injections) which are subject to an expected network SNR cut of 6, as lower SNRs are considered “hopelessly” undetectable LVK (2023, 2021b). We thus exclude samples at higher distance than any injection with comparable masses (similar data cleaning was performed in Sadiq et al. (2024b)).
@@ III.2 Selection effects
For the search sensitivity estimate or selection function, we employ an analytic approximant Lorenzo-Medina and Dent (2025) which accurately fits the probability of detection $p_{\mathrm{det}}$ at the 1 per year FAR threshold for injection (simulated signal) results released with the catalog LVK (2021b). The $p_{\mathrm{det}}$ approximant is a function of masses, effective inspiral spin $\chi_{\rm eff}$ Ajith et al. (2011) and luminosity distance, marginalizing over the sky direction, source orientation and spin orientations at given $\chi_{\rm eff}$. For our main results, we approximate $\chi_{\mathrm{eff}}$ to zero when evaluating the selection function for simplicity, motivated by findings in Abbott et al. (2023b) which support small effective spins for the observed BBH population.
@@
Figure 2: Detection probability $p_{\mathrm{det}}$ as a function of the primary mass $m_{1}$ and luminosity distance $d_{L}$ for parameter estimation samples from 69 BBH events in GWTC-3. The top panel assumes a power-law distribution for the mass ratio, $p(q)\propto q^{1.26}$, while the bottom panel uses posterior sample values of secondary mass $m_{2}$.
@@
The fitted detection probability is a function of the redshifted masses $(1+z)m_{1,2}$ which describe the GW signal at the detectors. To relate the redshift to $d_{L}$ we assume a standard flat $\Lambda$CDM cosmology with $H_{0}=67.9\,$km/s/Mpc and $\Omega_{M}=0.3065$ Ade et al. (2016), as also used for LVK parameter estimation; changes in the assumed cosmology, particularly in $H_{0}$, would have minor effects on the reconstructed source mass spectrum.
In some regions of the parameter space, principally at high distance, $p_{\mathrm{det}}$ can become extremely small. Our KDE of detected events will also take low values in regions with small numbers of samples. In principle, our estimate of the astrophysical merger density is proportional to the ratio of the KDE to $p_{\mathrm{det}}$; however, in the limit of very small $p_{\mathrm{det}}$ this estimate will become numerically unstable. To mitigate or regularize this issue, we limit the range of $p_{\mathrm{det}}$ used in the analysis by replacing (capping) values below a predefined threshold $p_{\mathrm{det}}=0.1$ by the threshold value, as demonstrated in the context of LISA mock data in Sadiq et al. (2024b). As shown in Fig. 2, most samples have $p_{\mathrm{det}}$ values above or around $0.1$, but a small fraction have significantly lower values, particularly at larger distances or for higher primary masses. As a result of this capping or regularization we expect the rate estimate to be biased downwards, and eventually to approach zero, in regions with very small $p_{\mathrm{det}}$.

### v2

# Looking To The Horizon: Probing Evolution in the Black Hole Spectrum With Gravitational Wave Catalogs

@@ Abstract
The population of black holes observed via gravitational waves currently covers the local universe up to a redshift $z\lesssim 1$, for the most massive merging binaries, or $z\lesssim 0.25$ for low-mass BH binaries (BBH). Evolution of the BBH mass spectrum over cosmic time will be a significant probe of formation channels and environments. We demonstrate a reconstruction of the BBH merger rate, allowing for general dependence on binary masses and luminosity distance or redshift and accounting for selection effects, via iterative kernel density estimation (KDE) with optimized multidimensional bandwidths. Performing such reconstructions under a range of detailed assumptions, we see no significant evidence for the evolution of BBH masses with redshift, over the range where detected events are available. At most, possible trends towards increasing merger rate with redshift for primary masses $m_{1}\gtrsim 50\,M_{\odot}$, or towards decreasing merger rate with redshift for primary masses $m_{1}\lesssim 40M\odot$ may be supported. We compare these findings with previous investigations and caution against over-interpreting the current, sparse, data. Significantly upgraded detectors and/or facilities, and longer observing times, are required to harness any correlations of the BBH mass distribution with redshift.
@@ I Introduction
One essential aspect of population analysis is understanding how the masses of BBHs correlate with redshift, which sheds light on their formation environments and underlying astrophysical processes. Recent studies have employed advanced statistical techniques to explore a possible correlation between BBH mass and redshift. In Rinaldi et al. (2024), a hierarchical Gaussian mixture is applied to changing properties of BBH over cosmic history, finding evidence of two distinct groups of black holes with different masses and redshifts, a lack of symmetric binary mergers, and a potential connection between black hole mass and the environment in which they formed. On the other hand Heinzel et al. (2024a) investigates the similar correlation using gravitational wave data with a method that avoids strong assumptions Heinzel et al. (2024b), finding limited evidence for evolution in black hole masses over cosmic time, but possible connections between spin, mass ratio, and distance may exist. Non-parametric population reconstruction has also been proposed to investigate more general questions in GW cosmology and astrophysics Ng et al. (2024); Fabbri et al. (2025).
@@
Most recently, in Lalleman et al. (2025) the authors focus on the power-law distribution and peak features in the primary mass, carrying out a parameterized change-point analysis to investigate possible changes in such features over redshift: they find no significant evidence of evolution, although various types of variation in the mass spectrum cannot be excluded.
@@
This study investigates how the mass distribution of BBHs may evolve with redshift using publicly available gravitational wave observations. We employ an improved version of a non-parametric method based on iterative kernel density estimation (KDE), building on previous studies Sadiq et al. (2022, 2024a, 2024b) in combination with an accurate fit of LVK search sensitivity Lorenzo-Medina and Dent (2025). Any evidence for such evolution would provide valuable insights into the astrophysical processes and formation channels that govern BBH systems. In Sadiq et al. (2022), a preliminary study of the density of detected BBH over mass and luminosity distance/redshift, using an earlier version of our adaptive KDE method but without accounting for selection effects, gave a qualitative indication that peak (local maximum) features in the mass distribution persisted over a range of distances. Here, we revisit that initial hint with several methodological improvements, including a self-consistent treatment of event measurement uncertainties Sadiq et al. (2024a) and, crucially, an accurate estimate of the selection function over binary masses and distance introduced in Lorenzo-Medina and Dent (2025) that enables us to probe the astrophysical rate density.
@@
A critical point given the current small statistics of GW detections is how far statistical inferences and scientific conclusions depend on model assumptions; and conversely, how sensitive they are to fluctuations in the observed GW event set, considered as a random sample. Clearly, little information can be obtained in regions of parameter space where no events are detected.
@@
The most obvious statistical limitation is the selection function restricting the visibility of BBH mergers at finite distance, due to detector and search sensitivity: the selection function is also strongly dependent on binary component masses (see e.g. discussion in Abbott et al. (2023a); Fishbach and Holz (2017), and Lorenzo-Medina and Dent (2025)). Thus, the BBH mass spectrum around $\sim\!10\,M_{\odot}$ can only be investigated out to fairly small redshift ($z\lesssim 0.25$) with current data. However even at higher masses, for which the current horizon of detectability extends farther, our knowledge of the BBH population is limited simply by the scarcity of mergers. We already detect a high fraction of high-mass BBH at low redshift, but this is still a very small number of events: thus, only significantly longer observing times will eventually enable an accurate local measurement.
Issues of selection bias and finite number statistics were explored in a study of a strongly redshift-dependent mock data population of “light seed” black holes observable by LISA, using the same KDE-based method, in Sadiq et al. (2024b). There, we found that the true population could be reconstructed by the method, but only within parameter regions with sufficient numbers of detections. For BBH currently observed by LIGO-Virgo, our knowledge of the selection function, as well as inspection of detected events, indicate that uncertainties in any current measurement of the redshift-dependent mass function are necessarily large. Thus, nontrivial model assumptions would probably be necessary to obtain any significant evidence of evolution.
@@
II Population reconstruction with iterative reweighted KDE
@@ II.1 Summary of previous work
Here $\hat{f}_{0}$ is an initial pilot density estimate obtained by setting $\lambda_{i}=1$ and $\alpha$ is the bandwidth sensitivity parameter (lying between 0 and $1$), $g$ being a normalization factor. The optimized choices of the kernel covariance (i.e. bandwidth) and $\alpha$ are determined by maximizing a likelihood figure of merit, evaluated via cross-validation Sadiq et al. (2022):
$\log\mathcal{L}_{\rm Kfold}=\sum_{k=1}^{K}\sum_{i\in\text{Fold}_{k}}\ln\hat{f}_{{\rm Fold}-k}(X_{i}).$ (3)
Here the data is partitioned into $K$ ‘folds’ of equal size; $\hat{f}_{{\rm Fold}-k}$ is the KDE trained on all folds except $\text{Fold}_{k}$. The log-likelihood score penalizes relative errors, making it suitable for estimation across a wide dynamic range of densities.
As we use detected events to optimize and evaluate the KDE, reconstruction of the actual astrophysical event density also requires an estimate of the selection function, as discussed below in Sec. III.2.
In Sadiq et al. (2024a), we improved the accuracy of our method by employing an iterative approach similar to Expectation-Maximization (EM) algorithms. The uncertainty in measured individual event properties, which for GW detections is often large, is represented by parameter estimation (PE) samples obtained by Bayesian inference (see e.g. Veitch et al. (2015)) using a standard default prior over mass and luminosity distance. Since these priors do not reflect the actual astrophysical population, the resulting parameter estimates are necessarily biased. Our iterative reweighting method is designed to correct for this bias self-consistently: we reweight the samples for each event at a given iterative step using density estimates from the previous step, adjusted for the selection function. We expect these estimates to, on average, have decreasing bias with successive iterations, as the KDE likewise more closely approaches the true (detected) distribution.
At every reweighting step, we draw Poisson (1) samples for each detected event Sadiq et al. (2024a), without replacement in the case of >1>1 samples, to avoid corner cases in cross-validation where the same sample point could appear in different folds. As with Markov Chain Monte Carlo methods, we desire to reach a steady state where successive iterations are fair draws from a distribution of density estimates representing the uncertainties on a reconstructed population: both due to finite event statistics and to individual event measurement errors.
We monitor the autocorrelation of the optimized bandwidth and adaptive parameter over iterations as diagnostics for the approach to stationarity. We discard an initial burn-in phase, and then establish a buffer of 100 iterations from which we compute a mean KDE; finally we obtain $500$ or more independent population estimates by bootstrap resampling as above, but using this buffer mean KDE to determine the reweighting of PE samples (in place of the previous Markov chain iteration). We quantify the population distribution via the median and 5th and 95th percentiles of these independent estimates.
This method is a computationally efficient approach to obtain self-consistent density and rate estimates, including bootstrap uncertainties. However, the method as previously applied is limited by the use of spherical (isotropic) kernels, implying $\Sigma=h^{2}\mathbb{I}$ in standardized coordinates, where $\mathbb{I}$ represents the identity matrix. This restriction can introduce bias in the estimates when the data follow distinct distributions, or have significantly different measurement uncertainties, over different dimensions.
II.2 Improvement to this work
Here we address this limitation by allowing for kernels with different relative bandwidths along each dimension, while the overall kernel scale varies between data points as already described for adaptive KDE. For a data point $\vec{X}_{i}$, the kernel contribution is proportional to
@@ II.2 Improvement to this work
This is not the most general linear data transformation, which would be $\vec{y}=H^{-1}\vec{x}$ where $H^{-1}$ is any invertible real matrix, corresponding to an off-diagonal kernel bandwidth matrix. We choose not to pursue this more complicated option: such kernels might increase apparent correlations between parameters, thus given our choice of diagonal kernel we expect our conclusions on correlations or evolution to be conservative.
The $D$-dimensional bandwidths $h_{1},\ldots h_{D}$ need to be optimized, along with the adaptive parameter $\alpha$. Given the dimensionality of this hyper-parameter space, instead of the grid search used in previous works we switch to a generalized optimization algorithm, in this case Nelder-Mead. Our cross-validated likelihood figure of merit is not invariant under rescaling coordinates: to ensure stable optimization we evaluate it over the $\vec{x}$ coordinates, i.e. without rescaling  [^1] Standardization also changes the likelihood figure of merit, but only by a constant for a given data set..
Figure 1: Verification of the multidimensional bandwidth optimization using mock data consisting of a mixture of two Gaussian components, represented by red ++ symbols: contours and blue shading show the KDE in each case. Top: Adaptive KDE with an isotropic kernel (in standardized coordinates): the single bandwidth parameter $h$ and adaptive parameter $\alpha$ optimize the cross-validated likelihood. Bottom: Adaptive KDE with different optimized bandwidths $h_{1,2}$ for each coordinate.
We validate the multi-dimensional kernel optimization using synthetic data comprising a bimodal Gaussian mixture. This dataset is useful for testing because the isotropic (spherical) kernel lacks the flexibility to capture its multimodality by using different bandwidths per parameter, even after standardizing the data.
We first apply adaptive KDE with an isotropic kernel in standardized coordinates, optimizing a single bandwidth parameter $h$ and adaptive parameter $\alpha$ to maximize the likelihood via leave-one-out cross-validation. However the resulting estimate, shown on the left of Fig. 1, is under-smoothed along one of the two dimensions, introducing excess variance and failing to represent the true distribution. Our improved KDE with independently optimized bandwidths $h_{1}$ and $h_{2}$, shown in Fig. 1 (right), better accommodates the anisotropic nature of the data, successfully recovering the bimodal Gaussian mixture, as well as significantly increasing the likelihood.
@@ III Application to GWTC-3 data
We apply the reweighted adaptive KDE with multi-dimensional bandwidth optimization to the results from parameter estimation LVK (2021a) of gravitational-wave observations cataloged in GWTC-3 Abbott et al. (2023a, c).
@@ III.1 Event selection, parameter estimation and data cleaning
Following the approach in our earlier studies Sadiq et al. (2022, 2024a), we selected 69 BBH candidate events from GWTC-3 with false alarm rates below 1 per year, excluding GW190814, an outlier event with an unusually low secondary mass ($m_{2}\simeq 2.6,M_{\odot}$, $q\simeq 0.11$), as its nature remains uncertain: either a very massive neutron star or a light black hole, which is inconsistent with the bulk BBH population we aim to analyze Abbott et al. (2020).
In LIGO-Virgo-KAGRA (LVK) catalog PE analyses, the prior over masses and $d_{L}$ is uniform over redshifted masses $m_{\{1,2\}z}=m_{\{1,2\}}(1+z)$ and over Euclidean volume Abbott et al. (2021a), thus is proportional to $(1+z)^{2}$ $d_{L}^{2}\mathrm{d}m_{1}\mathrm{d}m_{2}\mathrm{d}d_{L}$ (e.g. Callister (2021)). Samples reweighted to correspond to a “cosmological” prior uniform over comoving volume and source-frame time, though still uniform over $m_{z}$, are also available and are the basis of quoted event parameters Abbott et al. (2023a). We use the original samples and apply a factor proportional to the inverse prior $(1+z)^{-2}d_{L}^{-2}$ to our estimate of $p(m_{1}[,m_{2}],d_{L})$ in the reweighting step of our algorithm.
For two events with relatively low SNR, GW190719_215514 and GW190805_211137, we find a small number of PE samples at high distance which appear to represent the prior distribution, which is formally divergent towards large distance, rather than estimating the parameters of the detected merger. The expected (optimal) SNRs for these samples are unusually small. We compare the distances of PE samples with those of simulated signals (injections) which are subject to an expected network SNR cut of 6, as lower SNRs are considered “hopelessly” undetectable LVK (2023, 2021b). We exclude samples at higher distance than any injection with comparable masses (similar data cleaning was performed in Sadiq et al. (2024b)).
@@ III.2 Selection effects
For the search sensitivity estimate or selection function, we employ an analytic approximant Lorenzo-Medina and Dent (2025) which accurately fits the probability of detection $p_{\mathrm{det}}$ at the 1 per year FAR threshold for injection (simulated signal) results released with the catalog LVK (2021b). The $p_{\mathrm{det}}$ approximant is a function of masses, effective inspiral spin $\chi_{\rm eff}$ Ajith et al. (2011) and luminosity distance $d_{L}$, marginalizing over the sky direction, source orientation and spin orientations at given $\chi_{\rm eff}$. For our main results, we approximate $\chi_{\mathrm{eff}}$ to zero when evaluating the selection function for simplicity, motivated by findings in Abbott et al. (2023b) which support small effective spins for the observed BBH population.
@@
Figure 2: Detection probability $p_{\mathrm{det}}$ as a function of the primary mass $m_{1}$ and luminosity distance $d_{L}$ for PE samples from 69 BBH events in GWTC-3. The top panel assumes a power-law distribution for the mass ratio, $p(q)\propto q^{1.26}$, while the bottom panel uses posterior sample values of secondary mass $m_{2}$.
@@
[…]
In some regions of the parameter space, principally at high distance, $p_{\mathrm{det}}$ can become extremely small. Our KDE of detected events will also take low values in regions with small numbers of samples. In principle, our estimate of the astrophysical merger density is proportional to the ratio of the KDE to $p_{\mathrm{det}}$; however, in the limit of very small $p_{\mathrm{det}}$ this estimate will become numerically unstable. To mitigate or regularize this issue, we limit the range of $p_{\mathrm{det}}$ used in the analysis by replacing (capping) values below a predefined threshold $p_{\mathrm{det}}=0.1$ by the threshold value, as demonstrated in the context of LISA mock data in Sadiq et al. (2024b). As shown in Fig. 2, most samples have $p_{\mathrm{det}}$ values above or around $0.1$, but a small fraction have significantly lower values, particularly at larger distances or for higher primary masses. As a result of this capping or regularization we expect the rate estimate to be biased downwards, and eventually to approach zero, in regions with very small $p_{\mathrm{det}}$. In any case no useful information is available about the rate density in such “censored” regions, as they were dubbed in Rinaldi et al. (2024).